OpenStudy (anonymous):

y-int??? f(x)=3x^3-17x^2+18x+8

myininaya (myininaya):

y=8 is y intercept

myininaya (myininaya):

set x=0 and solve for y

myininaya (myininaya):

(0,8) y intercept

OpenStudy (anonymous):

3x^3-17x^2+18x+8=0

OpenStudy (anonymous):

how do u find the x-int??

myininaya (myininaya):

yes thats how yet set 3x^3-17x^2+18x+8=0 solve solve for x

myininaya (myininaya):

x=2 works

OpenStudy (anonymous):

i have stuck with x-int i dont know why i cant find it...

OpenStudy (anonymous):

(x-4)(x-2)(3x+1)

myininaya (myininaya):

2| 3 -17 18 8 | 6 -22 -8 ----------------- 3 -11 -4 | 0

OpenStudy (anonymous):

:)..how did u get x=2??? it's right but i wonder know why??

OpenStudy (anonymous):

you can guess using the rational root theorem... but there's a lot of guesses you can make for the root. It's somewhat trial and error unless you use calculus to determine the interval of one of the roots.

myininaya (myininaya):

3x^2-11x-4=0 \[x=\frac{11 \pm \sqrt{121-4(3)(-4)}}{6}=\frac{11 \pm \sqrt{121+48}}{6}=\frac{11 \pm \sqrt{169}}{6}=\frac{11 \pm 13}{6}\]

myininaya (myininaya):

right its a guessing thing

myininaya (myininaya):

i mean not really guessing you have a set of rational numbers and you can do trial and error i just chose 2 and it worked

OpenStudy (anonymous):

all the rational roots would be ±1, ±2, ±4, ±8, ±1/3, ±2/3, ±4/3, ±8/3.

myininaya (myininaya):

you look at factors of 8 and you look at factors or 3 and you also ook at factors of 8/3

myininaya (myininaya):

and he listed them all

myininaya (myininaya):

so we see we have two rational answers using the quad formula they should be in his list

myininaya (myininaya):

we have x=24/6=4 and x=-2/6=-1/3

OpenStudy (anonymous):

ok... thanks for everyone helping me!!!!